trains a and b are traveling in the same direction on parallel tracks. train a is traveling at 80 miles per hour and train b is traveling at 90 per hour train a passes a station at 5:20 am. if train b passes the same station at 5:50 am, what time will train b catch up to train a?
Answer by checkley77(12844)

80T=90(T-.5)
80T=90T-45
80T-90T=45
-10T=-45
T=-45/-10
T=4.5 HOUR AFTER a LEAVES THE STATION THEY WILL MEET. OR @ 9:50.
PROOF
80*4.5=90(4.5-.5)
360=90*4
360=360